OpenStudy (anonymous):

What is the standard form of the equation of a circle that has its center at (-2, -3) and passes through the point (-2, 0)? a) (x + 2)2 + (y + 3)2 = 9 b) (x − 2)2 + (y − 3)2 = 16 c) (x − 2)2 + (y − 3)2 = 4 d) (x + 2)2 + (y + 3)2 = 16 e) (x − 2)2 + (y + 3)2 = 9 please help !

OpenStudy (zzr0ck3r):

(x-a)^2+(y-b)^2 = r^2 is the equation of a circle with center (a,b) and radius r

OpenStudy (zzr0ck3r):

what is the radius? hint: the distance between two points (a,b) (c,d) is given by \[\sqrt{(a-c)^2+(b-d)^2}\]

OpenStudy (zzr0ck3r):

are you there @tracynicolexoxo

OpenStudy (anonymous):

yeah im here sorry lol

OpenStudy (zzr0ck3r):

do you understand?

OpenStudy (anonymous):

not really im horrible at math